# SurveyG: A Multi-Agent LLM Framework with Hierarchical Citation Graph for Automated Survey Generation

### v2

@@ 3.2.2. Full Paper Completion
LLM Model Coverage Structure Relevance Synthesis Critical Analysis
Claude AutoSurvey 73.6 64.5 80.2 51.1 45.8
SurveyForge 81.8 78.4 89.1 75.4 70.3
SurveyX 74.2 71.3 82.3 62.4 52.7
SurveyG 88.1 87.9 93.6 80.2 77.3
GPT AutoSurvey 90.7 86.2 89.3 84.6 82.3
SurveyForge 94.2 87.3 94.8 88.6 88.5
SurveyX 89.3 84.4 88.5 79.2 78.4
SurveyG 95.7 88.5 95.1 92.2 90.5
Deepseek AutoSurvey 86.5 80.4 87.8 77.3 72.4
SurveyForge 89.4 86.5 92.5 84.1 80.7
SurveyX 85.4 81.2 89.4 78.2 75.6
SurveyG 88.7 82.7 93.4 85.5 81.6
Gemini AutoSurvey 94.2 70.8 96.5 83.6 84.2
SurveyForge 94.9 93.1 98.7 93.5 94.2
SurveyX 90.2 71.2 95.6 82.8 84.5
SurveyG 96.2 89.4 97.6 93.6 94.9
Table 1. LLM-as-a-judge evaluation of generated surveys. Each LLM evaluates four generation models across content quality dimensions.
Model Recall Precision F1 Score
AutoSurvey 82.25 77.41 79.76
SurveyForge 88.34 75.92 81.66
SurveyX 85.23 78.12 81.52
SurveyG (ours) 90.60 76.32 83.49
Ground Truth 92.53 86.42 89.34
Table 2. Performance comparison of different models on citation quality metrics.
@@ 4.1.1. Baselines
We compare SurveyG with three state-of-the-art systems. AutoSurvey (Wang et al., [n. d.]) and SurveyX (Liang et al., 2025) represent multi-stage frameworks for automated survey generation, with SurveyX enhancing AutoSurvey through structured knowledge extraction and outline optimization. SurveyForge (Yan et al., 2025), in contrast, leverages human-written survey papers from related domains as prior knowledge for heuristic outline generation, guided by a memory-driven scholar navigation agent that retrieves high-quality references for composing new surveys.
4.1.2. Dataset Evaluation
To assess the generalizability and robustness of SurveyG, we evaluate it on ten diverse computer science topics from the SurGE benchmark (Su et al., 2025), which includes 205 ground-truth surveys and over one million papers. We recruited 20 domain experts, CS Ph.D. students from QS 5-star universities and senior AI research engineers, to curate high-quality reference papers and select one representative ground-truth survey per topic. The same experts also served as human evaluators, assessing the coherence, coverage, and factual accuracy of generated surveys. Additional details on ground-truth construction and evaluation protocols are provided in Appendix B.
@@ 4.1.3. Implementation Details
For fair comparison, we strictly follow the experimental settings of prior works (Yan et al., 2025; Wang et al., [n. d.]; Liang et al., 2025). Specifically, we retrieve 1,500 candidate papers for outline generation and 60 relevant papers for each chapter-writing stage, identical to SurveyForge (Yan et al., 2025). All experiments use GPT-4o-mini-2024-07-18 as the backbone model for both Writing and Evaluation Agents, consistent with previous studies. We generate surveys for ten predefined topics, each with ten independent trials (100 surveys in total), and report averaged results for stability. To assess cross-model robustness, we additionally test Gemini-2.5-Flash. For evaluation, advanced LLM judges: GPT-4o-2024-08-06, Claude-3.5-Sonnet-20241022, DeepSeek-V3.2-Exp, and Gemini-2.5-Pro are used to score both outlines and full survey texts. We set the iteration number $T_{MAX}$ to 2, the same as AutoSurvey and SurveyForge.
@@ 4.1.4. Evaluation Metrics
We evaluate the generated outputs along three dimensions: outline, content, and citation quality. For Outline Quality, we follow the evaluation protocol of (Yan et al., 2025), using the same prompt. For content quality, we adopt the five widely used metrics from prior benchmarks (Wang et al., [n. d.]; Liang et al., 2025; Yan et al., 2025; Su et al., 2025): Coverage, Structure, Relevance, Synthesis, and Critical Analysis. Six metrics are rated on a 0-100 scale by both LLM and human judges, measuring completeness, organization, topical alignment, integrative reasoning, and analytical depth. For citation quality (Wang et al., [n. d.]; Gao et al., 2023), we evaluate the factual consistency between claims and their cited references using a Natural Language Inference model, reporting Citation Recall (ratio of supported claims), Citation Precision (ratio of valid references), and Citation F1 as their harmonic mean. Details of these metrics are provided in Appendix B.
Model Score Win Rate Comparative Win Rate Human Eval Overall Score
SurveyForge 38.85% 27.75% 36.00% 82.55%
SurveyG (ours) 61.15% 72.25% 64.00% 87.67%
Table 3. Comparison of models across full paper evaluation metrics.
Model Score Win Rate Comparative Win Rate Human Eval Overall Score
SurveyForge 45.00% 42.00% 45.00% 90.00%
SurveyG (ours) 55.00% 58.00% 55.00% 95.00%
Table 4. Comparison of models based on outline evaluation metrics.
4.2. Evaluation on Content Quality
As shown in Table 1, SurveyG consistently achieves the highest scores across nearly all metrics and evaluation models, demonstrating strong generalization and robustness under different LLM judges. In particular, SurveyG shows notable gains in Synthesis and Critical Analysis, reflecting its ability to integrate information and identify research gaps through the use of a hierarchical citation graph and multi-level summarization prompts. SurveyForge ranks second overall, outperforming SurveyX and AutoSurvey in Coverage and Structure due to its heuristic use of human-written surveys as prior knowledge. However, it remains less effective than SurveyG, which achieves superior organization and analytical depth without relying on human-written inputs, instead leveraging structured summarization and cross-community reasoning within the hierarchical framework.
4.3. Evaluation on Ground Truth
@@
4.4. Evaluation on Citation Quality
The experimental results presented in Table 2 clearly indicate that the SurveyG (ours) model sets a new standard for automated survey generation, demonstrating superior citation quality compared to existing systems. SurveyG achieves the highest Recall at 90.60 and the best F1 Score at 83.49. This high Recall figure is particularly notable, as it is very close to the Ground Truth Recall of 92.53, suggesting SurveyG is highly effective at comprehensively identifying and linking relevant literature. While SurveyX holds the lead in Precision (78.12), SurveyG’s significantly better F1 Score confirms its overall advantage in balancing the inclusion of necessary citations with the exclusion of irrelevant ones. In summary, SurveyG (ours) surpasses AutoSurvey, SurveyForge, and SurveyX in overall performance, demonstrating a marked improvement in the reliability and comprehensiveness of citations in generated surveys.
4.5. Human Evaluation
To validate our automated evaluation framework, we conducted a comparative assessment between SurveyG (ours) and SurveyForge across ten topics. We adopted a win-rate–based evaluation protocol, presenting anonymized outputs from both systems to domain experts and the automated evaluation system. Human experts were selected based on topic relevance and possessed extensive research experience in their respective domains. As summarized in Table 3, three complementary metrics were used to quantify performance differences. The Score Win Rate measures how often a model receives a higher absolute score from the LLM evaluator. The Comparative Win Rate reflects the frequency with which the LLM selects a model’s paper as superior in pairwise comparisons. The Human Evaluation Win Rate represents the proportion of times human experts preferred outputs from one model over the other. Under this framework, SurveyG consistently outperforms SurveyForge, achieving a Score Win Rate of $61.15\%$, a Comparative Win Rate of $72.25\%$, a Human Evaluation Win Rate of $64.00\%$, and an overall score of $87.67\%$. In contrast, SurveyForge records $38.85\%$, $27.75\%$, $36.00\%$, and $82.55\%$, respectively. These results demonstrate strong consistency between automated and human evaluations, confirming that our framework reliably captures expert-level judgment while maintaining scalability and efficiency.
As shown in Table 4, three complementary metrics were adopted to evaluate model performance on outline generation. Under this evaluation framework, SurveyG (ours) consistently outperforms SurveyForge, achieving a Score Win Rate of $55.00\%$,a Comparative Win Rate of $58.00\%$, a Human Evaluation score of $55.00\%$, and an Overall Score of $95.00\%$. In contrast, SurveyForge records $45.00\%$, $42.00\%$, $45.00\%$, and $90.00\%$, respectively. These findings demonstrate that our model produces higher-quality outlines, confirming that constructing a Hierarchical Citation Graph for paper retrieval in outline generation is a principled and effective approach.
4.5.1. Details of Human Evaluation
@@
Table 5. Inter-rater agreement between LLM and human evaluations
4.6. Cost estimation
The SurveyG framework generates survey papers with an average length of approximately 64k tokens, comparable to expert-written surveys. Each subsection is produced using around 12k input tokens and 800 output tokens. Additionally, the Evaluation Agent in the RAG loop performs one assessment per subsection, consuming approximately 3.7k input and 700 output tokens. Using this configuration, the total cost for generating a full 64k-token survey is estimated at \$1.5-\$1.7, depending on model pricing and API parameters. These results highlight the cost-effectiveness and scalability of the SurveyG framework for producing high-quality, large-scale literature surveys.
@@ 4.7. Ablation Studies
In Table 6, to evaluate the contribution of each component in our architecture, we present ablation study results comparing our full model against variants with specific components removed. Our full version achieves the best overall performance, demonstrating the effectiveness of our complete architecture. The inclusion of RAG significantly enhances Coverage (91.98) and Relevance (94.81) compared to the w/o MA variant (89.26 and 91.50), as it supplements the model with additional contextual information. More importantly, the full model outperforms both the w/o Vertical Traversal and w/o Horizontal Clustering variants, particularly in Structure scores (86.78), indicating that having all components working together enables superior information synthesis. This complete architecture allows the model to effectively integrate and organize information from multiple sources, resulting in more coherent and well-structured outputs across all evaluation metrics.
Variant Cov. Str. Rel. Syn. C.A
Full 91.98 86.78 94.81 85.44 83.34
w/o Vertical Traversal 90.51 85.05 93.91 84.92 84.28
w/o Horizontal Clustering 91.80 86.22 94.61 83.91 86.01
w/o MA 89.26 84.36 91.50 83.73 81.43
Table 6. We test three variants: (1) w/o Vertical Traversal uses only horizontal clustering and summarization within each layer; (2) w/o Horizontal Clustering performs only vertical path traversal from foundation papers; (3) w/o MA removes the Multi-Agent component.
@@ 5. Conclusion
In this work, we introduced SurveyG, an automated framework for survey generation that leverages hierarchical knowledge representation and multi-agent collaboration to address the limitations of existing LLM-based approaches. By modeling papers through a three-layer citation-similarity graph and employing both horizontal and vertical traversal strategies, SurveyG captures the structural relationships and evolutionary progress of research, enabling the creation of coherent and well-structured outlines. Through extensive evaluations, we demonstrated the effectiveness of our framework across diverse computer science topics. On the SurGE benchmark for autonomous computer science survey generation, both LLM-as-a-judge evaluations and human expert assessments demonstrate that SurveyG outperforms state-of-the-art frameworks across multiple dimensions.
@@ Appendix A Details about Traversal on Graph
Appendix B Experimental Setting
B.1. Survey Topics
We compiled a collection of ten representative survey papers covering diverse research areas, as summarized in Table 7. Each topic reflects an active line of inquiry within machine learning and natural language processing, providing a strong foundation for evaluating literature review generation.

### v3

@@ 3.1.1. Baselines
We compare SurveyG with three state-of-the-art systems: AutoSurvey (Wang et al., ), SurveyX (Liang et al., 2025), and SurveyForge (Yan et al., 2025). Details on each baseline are provided in Appendix C.1.
3.1.2. Dataset and Ground Truth Construction
We evaluate SurveyG on ten diverse computer science topics from the SurGE benchmark (Su et al., 2025), which comprises 205 human-authored surveys and over one million papers. To establish high-quality ground truth, we recruited 20 domain experts: CS Ph.D. students from QS 5-star universities with 3+ years of research experience and senior AI research engineers with publication records in top-tier venues.
The ground-truth selection followed a rigorous three-stage protocol: (1) Topic Selection: Ten representative research areas covering diverse subfields; (2) Survey Selection: For each topic, 2-3 domain-matched experts independently reviewed available surveys based on four criteria: comprehensive coverage, structural clarity, recency (within 5 years), and citation impact, with disagreements resolved through consensus; (3) Reference Curation: Experts identified seminal works, key methodological contributions, and recent advances essential for comprehensive coverage. The same experts served as human evaluators, assessing coherence, coverage, and factual accuracy. Complete details on topic descriptions, selection criteria, scoring rubrics, and inter-annotator agreement are in Appendix C.
@@ 3.1.3. Implementation Details
We adopt the experimental setup from prior studies (Yan et al., 2025; Wang et al., ; Liang et al., 2025). Specifically, we retrieve 1,500 candidate papers for outline construction, select 60 papers per chapter, and include 300 papers in each final survey. All methods, including SurveyG and the baselines, use GPT-4o-mini-2024-07-18 as the main backbone model and share the same crawled paper database released with SurveyForge. To assess cross-model robustness, we additionally test Gemini-2.5-Flash. For evaluation, we generate ten independent runs per topic (100 surveys in total). Four state-of-the-art LLMs: GPT-4o-2024-08-06, Claude-3.5-Sonnet-20241022, DeepSeek-V3.2-Exp, and Gemini-2.5-Pro, are used as judges. We set $T_{\text{MAX}}=3$ and DeBERTa-v3-large as our main text encoder model.
Evaluation Type Model Score Win Rate Comparative Win Rate Human Eval
Full Paper SurveyForge 38.85% 27.75% 36.00%
SurveyG (ours) 61.15% 72.25% 64.00%
Outline SurveyForge 45.00% 42.00% 45.00%
SurveyG (ours) 55.00% 58.00% 55.00%
Table 2. Win-rate comparison of SurveyForge and SurveyG on full-paper and outline evaluations. Score Win Rate indicates higher absolute evaluation scores, Comparative Win Rate reflects pairwise LLM preferences, and Human Eval denotes human judgment win rate.
@@ 3.1.4. Evaluation Metrics
We evaluate three dimensions: (1) Outline Quality following (Yan et al., 2025); (2) Content Quality using five metrics (Coverage, Structure, Relevance, Synthesis, Critical Analysis) rated 0-100 by LLM and human judges; (3) Citation Quality measuring factual consistency via NLI, reporting Citation Recall, Precision, and F1. Complete metric definitions and evaluation protocols are in Appendix C.3.
3.2. LLMs-as-a-judge evaluation
3.2.1. Evaluation on Content Quality
As shown in Table 1, SurveyG consistently achieves the highest scores across nearly all metrics and evaluation models, demonstrating strong generalization and robustness. SurveyG maintains low standard deviations (1.3-3.7 points) compared to baselines (1.1-5.4 points), indicating stable performance across topics and judges. Cross-judge analysis reveals remarkable consistency: SurveyG achieves top-2 performance under all four judges and ranks first in 16 out of 20 metric-judge combinations (80%). In contrast, SurveyForge shows variable performance, ranking first under Gemini but dropping to third under Claude’s stricter assessment. The particularly great improvements in Synthesis (+2.9 to +4.6 points over SurveyForge) and Critical Analysis (+1.9 to +7.0 points) reflect the framework’s ability to integrate information and identify research gaps through structured reasoning, achieving superior organization and analytical depth without requiring human-written survey inputs.
3.2.2. Evaluation on Ground Truth
@@
3.2.3. Evaluation on Citation Quality
[…]
Figure 6. Ablation analysis of the number of evaluation loop $T_{MAX}$.
Figure 7. Ablation analysis of the number of foundation papers $K$.
Model Recall Precision F1 Score
AutoSurvey 82.25(± 2.8)82.25\,($\pm\,$2.8) 77.41(± 3.2)77.41\,($\pm\,$3.2) 79.76(± 2.5)79.76\,($\pm\,$2.5)
SurveyForge 88.34(± 2.1)88.34\,($\pm\,$2.1) 75.92(± 3.5)75.92\,($\pm\,$3.5) 81.66(± 2.3)81.66\,($\pm\,$2.3)
SurveyX 85.23(± 2.6)85.23\,($\pm\,$2.6) $\mathbf{78.12}$(± 2.9)($\pm\,$2.9) 81.52(± 2.4)81.52\,($\pm\,$2.4)
SurveyG (ours) $\mathbf{91.40}$(± 1.8)($\pm\,$1.8) 77.83(± 2.7)77.83\,($\pm\,$2.7) $\mathbf{83.49}$(± 2.0)($\pm\,$2.0)
Ground Truth 92.53(± 1.5)92.53\,($\pm\,$1.5) 86.42(± 2.1)86.42\,($\pm\,$2.1) 89.34(± 1.7)89.34\,($\pm\,$1.7)
Table 3. Citation quality comparison across models.
@@ 3.2.3. Evaluation on Citation Quality
Table 4. Inter-rater agreement between LLMs and humans
3.3. Human Evaluation
To validate our evaluation framework, we conducted win-rate comparisons between SurveyG and the strongest baseline SurveyForge across ten topics, using three metrics: Score Win Rate (higher absolute scores), Comparative Win Rate (pairwise preferences), and Human Eval (expert judgments). As shown in Table 2, SurveyG substantially outperforms SurveyForge in full-paper generation (61.15% vs. 38.85% score win rate, 72.25% vs. 27.75% comparative win rate, 64.00% vs. 36.00% human preference). The particularly strong comparative win rate (72.25%) indicates that judges overwhelmingly prefer SurveyG in direct side-by-side comparisons. For outline generation, SurveyG maintains consistency, demonstrating that the hierarchical citation graph provides effective structured guidance from the earliest generation stage. The strong alignment between automated and human judgments (within 3-8%) validates our LLM-based evaluation framework’s reliability while maintaining scalability.
3.3.1. Details of Human Evaluation
Variant Cov. Str. Rel. Syn. C.A
Full 95.7 88.5 95.1 92.2 90.5
w/o Vertical Traversal 89.8 84.8 93.9 84.8 84.6
w/o Horizontal Clustering 91.5 85.2 93.6 83.7 86.8
w/o MA 93.2 85.6 91.2 88.7 87.2
Table 5. Ablation study on key components: (1) w/o Vertical Traversal: horizontal clustering only; (2) w/o Horizontal Clustering: vertical path traversal only; (3) w/o MA: no multi-agent refinement. Evaluated by GPT-4o.
@@ 3.4. Ablation Studies
Table 5 presents ablation results comparing our full model against variants with specific components removed. The full model achieves the best performance across all metrics, demonstrating that each component contributes meaningfully to survey quality. Removing vertical traversal causes the most significant degradation, particularly in Synthesis (-7.4 points) and Critical Analysis (-5.9 points), indicating that cross-layer exploration from foundational to frontier papers is crucial for integrative reasoning and identifying research gaps. The w/o horizontal clustering variant also shows substantial drops in Synthesis (-8.5 points), confirming that within-layer grouping of thematically related papers enables effective information organization. Removing the multi-agent refinement loop (w/o MA) decreases Relevance (-3.9 points) and Structure (-2.9 points), showing that iterative evaluation-agent feedback improves content focus and coherence. These results validate that the hierarchical citation graph’s bidirectional traversal mechanisms and multi-agent refinement work synergistically to produce comprehensive, well-structured surveys.
Ablation on number of $T_{\text{MAX}}$. We investigate the impact of evaluation iterations $T_{\text{MAX}}$ on generation quality (Figure 6). For outline generation, performance improves from mean = 3.5 (std = 0.6) at $T_{\text{MAX}}=1$ to mean = 4.8 (std = 0.30) at $T_{\text{MAX}}=3$, where iterative feedback effectively addresses structural gaps. Similarly, section content quality increases from 82.5 (std = 3.2) at $T_{\text{MAX}}=1$ to 90.6 (std = 2.4) at $T_{\text{MAX}}=3$, achieving full model performance (Structure = 88.5 $\pm$ 2.2, Synthesis = 92.2 $\pm$ 2.3, Critical Analysis = 91.2 $\pm$ 2.6). Beyond $T_{\text{MAX}}=3$, improvements become marginal (¡ 0.2 points) with increased variance, indicating diminishing returns. We therefore set $T_{\text{MAX}}=3$ as default, balancing quality, stability, and computational efficiency.
Ablation on number of $K$. We examine how the number of foundation papers $K$ affects vertical traversal across Structure, Synthesis, and Critical Analysis metrics (Figure 7). Performance is suboptimal at $K=6$ (scores = 80.9-81.8, std = 3.5-4.0) due to insufficient foundational coverage, improves at $K=8$ (scores = 84.0-85.1), and peaks at $K=10$ (scores = 84.6-85.9, std = 2.6-2.8). Synthesis shows the highest sensitivity to $K$ (5.0-point range), as comprehensive foundation-to-frontier paths are critical for cross-paper integration. Optimal $K$ varies by topic breadth: broader fields benefit from $K=10$, while focused domains achieve comparable results with $K=8$. Beyond $K=10$, performance plateaus or declines with increased variance. We recommend $K\in[8,10]$ based on topic characteristics.
Cost Estimation. The SurveyG framework generates survey papers with an average length of approximately 64k tokens, comparable to expert-written surveys. Each subsection is produced using around 12k input tokens and 800 output tokens. The Evaluation Agent in the RAG loop performs one assessment per subsection, consuming approximately 3.7k input and 700 output tokens. Using GPT-4o-mini pricing (\$0.150/1M input tokens, \$0.600/1M output tokens), the total LLM cost for generating a full 64k-token survey is \$1.5-\$1.7, depending on the number of sections and revision iterations. We provide details about computational efficiency in Appendix E.
More ablation studies. We compare SurveyG and SurveyForge under 2 different backbones: GPT-4o-mini and Gemni-2.5-Flash in Appendix B.
4. Related works
4.1. Long-form Text Generation
LLMs have achieved remarkable progress, yet generating long-form, coherent, and logically structured documents remains a persistent challenge (Bai et al., 2023; Dong et al., 2023; Han et al., 2024). Recent works have explored different strategies to address the long-context problem. For example, Chain-of-Agents (Zhang et al., 2024) introduces a multi-agent collaboration framework where worker agents process segmented portions of text and a manager agent synthesizes them into coherent outputs, alleviating focus issues in long contexts. LongAlign (Bai et al., ) proposes a recipe for long context alignment, combining instruction data construction, efficient batching, yielding strong gains on queries up to 100k tokens. Complementary to these, Xu et al. (Xu et al., 2023) systematically examine the trade-offs between retrieval-augmentation and context-window extension, showing that hybrid approaches can outperform both strategies alone. However, existing approaches often rely on raw reference texts, leading to inefficient retrieval, limited context utilization, and poor structural coherence in survey-like outputs.
4.2. Automatic Survey Generation
The automatic generation of literature reviews has been studied for over a decade, starting with multi-document summarization techniques that produced unstructured related work sections. Early systems, such as IBM Science Summarizer (Erera et al., 2019), focused on summarizing scientific articles, while more recent LLM-based methods like ChatCite (Li et al., 2024) and Susnjak et al. (Susnjak et al., 2025)’s domain-specific fine-tuning advanced the generation of comparative and knowledge-enriched reviews. Despite these advances, such methods primarily tackle summarization rather than the creation of fully structured survey papers. More recent systems, including AutoSurvey (Wang et al., ), InteractiveSurvey (Wen et al., 2025), SurveyForge (Yan et al., 2025), and SurveyX (Liang et al., 2025), propose end-to-end pipelines integrating RAG, clustering, or multi-agent strategies to automate survey construction. These methods improve structural coherence and formatting consistency while scaling to long-form survey content. Nevertheless, most frameworks still restrict users to fixed input-output modes, overlooking relationships among papers and limiting interactivity, which often results in surveys that lack flexibility, relational awareness, and depth.
5. Limitations and Ethical Considerations
Limitations.
SurveyG requires substantial computational resources for graph construction and multi-agent refinement, which may limit accessibility for researchers with constrained budgets. The framework’s effectiveness depends on the availability and quality of crawled papers, with access limitations to paywalled content potentially creating gaps in coverage. Our evaluation focuses exclusively on English-language computer science publications; generalization to other disciplines or languages remains unexplored.
Ethical Considerations.
Automated survey generation raises concerns about attribution and academic integrity. Users must critically review outputs to ensure proper credit and avoid inadvertent plagiarism. The citation graph inherently reflects existing biases in academic publishing, including systematic underrepresentation of work from marginalized researchers, non-prestigious institutions, and non-English publications.
@@ 6. Conclusion
In this work, we introduced SurveyG, an automated framework for survey generation that leverages hierarchical knowledge representation and multi-agent collaboration to address the limitations of existing LLM-based approaches. By modeling papers through a three-layer citation-similarity graph and employing both horizontal and vertical traversal strategies, SurveyG captures the structural relationships and evolutionary progress of research, enabling the creation of coherent and well-structured outlines. On the SurGE benchmark for autonomous computer science survey generation, both LLM-as-a-judge evaluations and human expert assessments demonstrate that SurveyG outperforms state-of-the-art frameworks across multiple dimensions.